# Commutator calculus and symbolic differentiation of matrix functions

### v2

@@ Abstract.
We propose a functional calculus which allows one to apply functions to the matrix anti-commutator/commutator operator. The calculus is introduced in a straightforward manner if the operators act on symmetric matrices, and it leads to a coordinate-free version of Daleckiĭ–Kreĭn formula. In this sense, the proposed calculus provides symbolic formulae for the derivatives of matrix-valued functions that are explicit and easy to use. We discuss several applications of the newly introduced calculus in continuum mechanics (Hencky logarithmic strain, objective rates, spin tensors, viscoelastic fluids) and in the theory of partial differential equations.
@@ 1. Motivation
for any $\mathbf{X}\in\mathbb{R}^{d\times d}$. These four operators represent the left/right matrix multiplication by $\mathbf{G}$ and the (halved) anti-commutator/commutator operators, respectively. For simplicity, in this work the matrix $\mathbf{G}$ is usually symmetric, denoted as $\mathbf{G}\in\mathbb{R}^{d\times d}_{\rm sym}$, although most of our conclusions also extend to the asymmetric case. Note that the operators $\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}},\operatorname{\mathsf{a\!c}}_{\mathbf{G}},\operatorname{\mathsf{a\!d}}_{\mathbf{G}}$ defined by (1.1) can be seen as fourth-order tensors whose components derive from the underlying matrix $\mathbf{G}$ and its spectral decomposition, which is the only place where the choice of coordinates matters. Otherwise, we shall adhere to the corresponding coordinate-free operator representations, taking a general functional-analytic approach. The need to apply a function to the operators in (1.1) stems from the problem of finding symbolic matrix derivatives of classical matrix functions such as exponential or logarithm. For instance, recall that a classical Lie group formula (cf. [14, Th. 5.4] and [34, Sect. 1.2., Th. 5]) reads as
@@
The expressions (1.2) and (1.4) show that it is desirable to look for other ways of interpreting the symbol $f(\operatorname{\mathsf{a\!d}}_{\mathbf{G}})$, where $f$ is not necessarily an entire function. Although one can interpret $f(\operatorname{\mathsf{a\!d}}_{\mathbf{G}})$ using general tools of functional analysis applied to the (self-adjoint) operator $\operatorname{\mathsf{a\!d}}_{\mathbf{G}}$, it turns out that in this case there exists an elegant approach that is particularly easy to work with and derives explicitly from the spectral decomposition of $\mathbf{G}$. Moreover, the functional calculus developed in this way not only gives meaning to expressions of type $f(\operatorname{\mathsf{a\!d}}_{\mathbf{G}})$, but also provides a tool for symbolic differentiation of matrix-valued functions due to an intrinsic connection with the Daleckiĭ–Kreĭn formula. For example, the Fréchet derivative of a matrix function $f$ can be conveniently expressed in terms of the matrix left and right multiplication operators as
@@ 2. Functions applied to the commutator
In this section, we introduce the calculus, see Definition 2.1, and study its elementary properties, see Theorem 2.2. Next we exploit the Cayley–Hamilton theorem and study the calculus in a low-dimensional setting, which is of interest in applications. For the sake of smooth text flow, the proofs are postponed to Section 5.
@@
Hereafter, we use the brackets $[m_{ij}]$ to denote a square matrix of size $d\times d$ with elements $m_{ij}$, $i,j=1,\ldots,d$. Using the spectral decomposition of $\mathbf{G}$, we now give an elementary definition of a function applied to any of the operators introduced in (1.1).
@@
Identity (2.2) defines a linear operator $f(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})$ on $\mathbb{R}^{d\times d}$, i.e., $f(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})\in\mathcal{L}(\mathbb{R}^{d\times d})$, and we often suppress the argument $\mathbf{X}$ when possible. The above definition covers functions of the commutator: If the function $h$ is defined in a neighborhood of the points $\frac{g_{i}-g_{j}}{2}$, $i,j=1,\ldots,d$, (which are the eigenvalues of $\operatorname{\mathsf{a\!d}}_{\mathbf{G}}$, cf. [34, Sect. 1.2., Lemma 8]), then
@@
The purpose of identification (2.3) is only to treat seamlessly functions with removable discontinuities, such as $\eta$ and $1/\eta$, avoiding the conditional definitions of these functions. In other words, we replace the functions with removable discontinuities by their continuous versions as we have no use of the former. We shall see later that this convention also allows for a very concise representation of matrix derivatives, where a special case of (2.2) is known as the Daleckiĭ–Kreĭn formula, see [11] and [24, Sect. 4.2]. In this respect, the merit of Definition 2.1 and of this work in general is that one can apply virtually the same formula to define the functions of the commutator as for the matrix derivatives, leading eventually to a comprehensive and elegant calculus that avoids the problems with convergence of power series.
@@ 2.1. Rules of the commutator calculus
    (Addition & multiplication) There holds
    $(f_{1}+f_{2})(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})=f_{1}(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})+f_{2}(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})$
@@
    $(f_{1}f_{2})(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})=f_{1}(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})f_{2}(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}}).$
@@ 2.2. Explicit formulae of function of the commutator in two and three dimensions
Let $\mathbf{G},\mathbf{g}\in\mathbb{R}^{d\times d}_{\rm sym}$ and $\{g_{i}\}_{i=1}^{d}$, be as in (2.1). Let $f$ be defined in a neighbourhood of the set $\{0,\frac{g_{i}-g_{j}}{2};i\neq j\}$ and consider its decomposition $f=f_{0}+f_{\rm odd}+f_{\rm even}$, where
@@
The right-hand sides of (2.12) and (2.13) can be, of course, written again only in terms of $\operatorname{\mathsf{a\!d}}_{\mathbf{G}}$, $\operatorname{\mathsf{a\!c}}_{\mathbf{G}}$ and their compositions; however, this does not simplify the result. We recall that the eigenvalues of any square matrix can be seen as certain functions of the principal invariants of that matrix. For example, if $d=2$, then one has
$g_{1,2}=\frac{1}{2}\operatorname{tr}\mathbf{G}\pm\frac{1}{2}\sqrt{\operatorname{tr}^{2}\mathbf{G}-4\det\mathbf{G}},$
which can be used to rewrite (2.12) as a function of $\mathbf{G}$ and $\mathbf{X}$ only, avoiding the spectral decomposition altogether. This hints at the fact that, using a more general, and likely more technical, approach, one could define $f(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})$ also for nonsymmetric square matrices $\mathbf{G}$. In fact, it seems that formulae (2.12) and (2.13) continue to hold in that case, although this is not proved here. We shall follow up on this observation in the next section (cf. Conjecture 3.7 below), where we address the derivatives of the matrix functions.
@@ 3. Symbolic differentiation of matrix functions
In this work, by matrix functions we mean those that are defined by the spectral decomposition as in (2.10). For matrix derivatives, there are several possible notions. The following theorem puts forward a unifying approach.
@@
[…]
@@
[…]
@@
    If $\mathbf{G}\in\mathbb{R}^{d\times d}_{\rm sym}$ and $f\in\mathcal{C}^{1}(I;\mathbb{R})$ for some interval $I\subset\mathbb{R}$ containing the spectrum of $\mathbf{G}$, then $\frac{df(\mathbf{G})}{d\mathbf{G}}$ fulfils the Daleckiĭ–Kreĭn [11] formula
@@
[…]
@@
which retrieves the well-known chain rule for the matrix commutator, cf. [31, Lem. 1], see also [9, (1.9)] for general self-adjoint operators.
@@ 3.1. Formulae for derivatives of special matrix functions
Now we have all the tools ready to calculate a symbolic derivative of a particular function. We remark that the task of differentiating a general matrix function was already investigated in the 1950s, see [33] and subsequent developments in [31] and references therein. Here, we shall be predominantly interested in the exponential, logarithm, trigonometric, and power functions, as these are the most important in applications. Unlike in the scalar case, the same matrix derivative can always be written in multiple ways, which can be useful in various situations. The list of formulae in the following theorem includes those that are known in literature (see [7, X.4] in particular); however, we provide alternative (simple) proofs based on the formula (3.1) or (3.4), see Section 5.
@@
[…]
@@
Formula (3.8) for the derivative of the matrix exponential appears probably in every book on Lie algebras; see, for instance, [14, Th. 5.4] or [34, Sect. 1.2., Th. 5]. Its counterpart (3.9), which is in fact just a combination of (3.8) with the Hadamard lemma (cf. Lemma 3.4 below), is much less common, but sometimes appears in proofs of the Baker–Campbell–Hausdorff formula. Symmetrized identities (3.10), (3.11) can be found in [24, Sect. 4] and in [16, Th. 10.13], however encoded in a different language, using the Kronecker representation, which eventually relies on the interpretation of power series. The inverted versions (3.14)–(3.17) seem to be not stated in literature this explicitly and for an arbitrary positive definite matrix $\mathbf{A}$. It also appears that the potential strength of (3.17) in physics and PDE applications has not been accordingly utilized so far, see Section 4 below. The identity (3.7) is well known and is the same as (3.5) mentioned above. Formula (3.12) is the exact counterpart of (3.5), although it is unlikely to be found in the literature, since it is based on Definition 2.1. On the other hand, identity (3.13) can be directly interpreted in a usual way and it is indeed well known (cf. [16, (11.10)]), since it follows from the integral definition of the matrix logarithm (see [16, (11.1)]). The formula for the derivative of the matrix power (3.22) can be directly applied to matrix-valued partial differential equations with diffusion; see Section 4.2 below. A similar formula to (3.19) is derived in [8] to solve a certain matrix equation, which can also be directly treated using (3.4). Finally, regarding the formulae for the matrix hyperbolic and trigonometric functions, these are obviously direct consequences of the previous identities for the matrix exponential. It seems that the only relatable result in the literature is [16, Th. 12.4], which uses the Kronecker representation.
We shall demonstrate in Section 4 below that Theorem 3.2 is very useful in symbolic matrix computations that are then used to obtain theoretical results. On the other hand, the question of optimal numerical evaluation of the above derivatives and matrix functions in general is a very different and highly nontrivial problematic, to which there corresponds an enormous literature, see the overview [23] and references therein just for the computation of the matrix exponential. Let us at least provide the following worked-out example that should clarify how the above formulae can be evaluated concretely.
@@
from Theorem 3. Note again that the right-hand side is the composition of two fourth-order tensors: part $\operatorname{\mathsf{a\!c}}_{\mathbf{A}^{-1}}$ is expected to be there in the commutative case and the rest depends on the commutator. For simplicity, consider the case $d=2$.
Although we could apply Theorem 2 directly, we shall use a little trick which consists of writing
@@
$=\frac{1}{2}(\mathbf{A}^{-1}\mathbf{X}+\mathbf{X}\mathbf{A}^{-1})+\frac{\theta(\delta)}{8\delta}\mathbf{A}^{-1}\big{(}(\log\mathbf{A})^{2}\mathbf{X}-2(\log\mathbf{A})\mathbf{X}(\log\mathbf{A})+\mathbf{X}(\log\mathbf{A})^{2}\big{)}$
@@
This could be, of course, further manipulated using, e.g., the Caley–Hamilton theorem for $\mathbf{A}^{-1}$, however the core idea should already be clear. Note that the resulting formula (3.36) indeed only sees the matrix $\mathbf{A}$ and not its spectral decomposition.
@@
see, e.g., [34, p. 3, (b)]. Hence, applying also (2.9) gives

### v3

@@ Abstract.
We propose a functional calculus which allows one to apply functions to the matrix anti-commutator/commutator operator. The calculus is introduced in a straightforward manner if the operators act on symmetric matrices, and it leads to a coordinate-free version of Daleckiĭ–Kreĭn formula. In this sense, the proposed calculus provides symbolic formulae for the derivatives of matrix-valued functions that are explicit and easy to use. We discuss several applications of the newly introduced calculus in continuum mechanics (Hencky logarithmic strain, objective rates, spin tensors, and viscoelastic fluids) and in the theory of partial differential equations.
@@ 1. Motivation
for any $\mathbf{X}\in\mathbb{R}^{d\times d}$. These four operators represent the left/right matrix multiplication by $\mathbf{G}$ and the (halved) anti-commutator/commutator operators, respectively. For simplicity, in this work the matrix $\mathbf{G}$ is usually symmetric, denoted as $\mathbf{G}\in\mathbb{R}^{d\times d}_{\rm sym}$, although most of our conclusions also extend to the asymmetric case. Note that the operators $\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}},\operatorname{\mathsf{a\!c}}_{\mathbf{G}},\operatorname{\mathsf{a\!d}}_{\mathbf{G}}$ defined by (1.1) can be seen as fourth-order tensors whose components derive from the underlying matrix $\mathbf{G}$ and its spectral decomposition, which is the only place where the choice of coordinates matters. Otherwise, we shall adhere to the corresponding coordinate-free operator representations, taking a general functional-analytic approach. The need to apply a function to the operators in (1.1) stems from the problem of finding symbolic matrix derivatives of classical matrix functions such as exponential or logarithm. For instance, recall that a classical Lie group formula (cf. [14, Th. 5.4] and [35, Sect. 1.2., Th. 5]) reads as
@@
The expressions (1.2) and (1.4) show that it is desirable to look for other ways of interpreting the symbol $f(\operatorname{\mathsf{a\!d}}_{\mathbf{G}})$, where $f$ is not necessarily an entire function. Although one can interpret $f(\operatorname{\mathsf{a\!d}}_{\mathbf{G}})$ using general tools of functional analysis applied to the (self-adjoint) operator $\operatorname{\mathsf{a\!d}}_{\mathbf{G}}$, it turns out that in this case there exists an elegant approach that is particularly easy to work with and derives explicitly from the spectral decomposition of $\mathbf{G}$. Moreover, the functional calculus developed in this way not only gives meaning to expressions of type $f(\operatorname{\mathsf{a\!d}}_{\mathbf{G}})$, but also provides a tool for symbolic differentiation of matrix-valued functions due to a formal similarity with the Daleckiĭ–Kreĭn formula. For example, it turns out that the Fréchet derivative of a matrix function $f$ can be conveniently expressed in terms of the matrix left and right multiplication operators as
@@ 2. Functions applied to the commutator
In this section, we introduce the calculus, see Definition 2.1, and study its elementary properties.
@@
Hereafter, we use the brackets $m_{ij}]$ to denote a square matrix of size $d\times d$ with elements $m_{ij}$, $i,j=1,\ldots,d$. Using the spectral decomposition of $\mathbf{G}$, we now give an elementary definition of a function applied to any of the operators introduced in (1.1).
@@
Identity (2.2) defines a linear operator $f(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})$ on $\mathbb{R}^{d\times d}$, i.e., $f(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})\in\mathcal{L}(\mathbb{R}^{d\times d})$, and we often suppress the argument $\mathbf{X}$ when possible. The above definition covers functions of the commutator: If the function $h$ is defined in a neighborhood of the points $\frac{g_{i}-g_{j}}{2}$, $i,j=1,\ldots,d$, (which are the eigenvalues of $\operatorname{\mathsf{a\!d}}_{\mathbf{G}}$, cf. [35, Sect. 1.2., Lemma 8]), then
@@
The purpose of identification (2.3) is only to treat seamlessly functions with removable discontinuities, such as $\eta$ and $1/\eta$, avoiding the conditional definitions of these functions. In other words, we replace the functions with removable discontinuities by their continuous versions as we have no use of the former. We shall see later that this convention also allows for a very concise representation of matrix derivatives, where a special case of (2.2) is known as the Daleckiĭ–Kreĭn formula, see [11] and [25, Sect. 4.2]. In this respect, the merit of Definition 2.1 and of this work in general is that one can apply virtually the same formula to define the functions of the commutator as for the matrix derivatives, leading eventually to a comprehensive and elegant calculus that avoids the problems with convergence of power series.
@@ 2.1. Rules of the commutator calculus
    (Addition & multiplication) If $s,p:U\times U\to\mathbb{R}$ are such that
    $s(x,y)=f_{1}(x,y)+f_{2}(x,y)\quad\text{and}\quad p(x,y)=f_{1}(x,y)f_{2}(x,y)$
    for all $x,y\in U$, then
    $s(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})=f_{1}(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})+f_{2}(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})$
@@
    $p(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})=f_{1}(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})f_{2}(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}}).$
@@
Using Definition 2.1, the proof of the above theorem is straightforward and is deferred to Section 5, as well as most of the proofs in this work.
@@ 2.2. Explicit formulae of function of the commutator in two and three dimensions
Let $\mathbf{G},\mathbf{g}\in\mathbb{R}^{d\times d}_{\rm sym}$ and $\{g_{i}\}_{i=1}^{d}$, be as in (2.1). Let $f$ be defined in a neighborhood of the set $\{\frac{g_{i}-g_{j}}{2}\}_{i,j=1}^{d}$ and consider its decomposition $f=f_{0}+f_{\rm odd}+f_{\rm even}$, where
@@
The right-hand sides of (2.12) and (2.13) can be, of course, written again only in terms of $\operatorname{\mathsf{a\!d}}_{\mathbf{G}}$, $\operatorname{\mathsf{a\!c}}_{\mathbf{G}}$ and their compositions; however, this does not simplify the result.
In the next section, we shall see that the derivatives of the classical matrix functions lead to expressions involving $f(\operatorname{\mathsf{a\!d}}_{\mathbf{G}})$ for certain functions $f$. Theorem 2.3 can be used to transform these expressions into a usual form, without relying on Definiton 2.1. In Example 3.3 below, we demonstrate this procedure. In addition to that, avoiding the determination of the matrix $\mathbf{Q}$ can be crucial in those applications where $\mathbf{G}$ itself is a function of some other variables and one has to investigate properties of the mapping $\mathbf{G}\mapsto f(\operatorname{\mathsf{a\!d}}_{\mathbf{G}})$ such as continuity. Furthermore, the absence of eigenspaces of $\mathbf{G}$ in Theorem 2.3 indicates that, using a more general, and likely more technical, approach, one could define $f(\operatorname{\mathsf{L}}_{\mathbf{G}},\operatorname{\mathsf{R}}_{\mathbf{G}})$ also for non-symmetric square matrices $\mathbf{G}$. In fact, it seems that formulae (2.12) and (2.13) continue to hold in that case, although this is not proved here. We shall follow up on this observation in the next section (cf. Conjecture 3.7 below), where we address the derivatives of the matrix functions.
@@ 3. Symbolic differentiation of matrix functions
In this work, by matrix functions we mean those that can be defined by the spectral decomposition as in (2.10). For matrix derivatives, there are several possible notions. The following theorem puts forward a unifying approach.
@@
$\frac{df(\mathbf{G})}{d\mathbf{G}}:=\frac{1}{2}\int_{-1}^{1}f^{\prime}(\operatorname{\mathsf{a\!c}}_{\mathbf{G}}+\,s\operatorname{\mathsf{a\!d}}_{\mathbf{G}})ds,$ (3.1)
@@
$\frac{df(\mathbf{G})}{d\mathbf{G}}:=\int_{0}^{1}f^{\prime}\big((1-t)\operatorname{\mathsf{L}}_{\mathbf{G}}+\,t\operatorname{\mathsf{R}}_{\mathbf{G}}\big)dt,$ (3.2)
@@
    If $\mathbf{G}\in\mathbb{R}^{d\times d}_{\rm sym}$ and $f\in\mathcal{C}^{1}(I;\mathbb{R})$ for some interval $I\subset\mathbb{R}$ containing the spectrum of $\mathbf{G}$, then $\frac{df(\mathbf{G})}{d\mathbf{G}}$ fulfills the Daleckiĭ–Kreĭn [11] formula
@@
$\frac{de^{\mathbf{G}}}{d\mathbf{G}}\mathbf{X}=\int_{0}^{1}e^{(1-t)\mathbf{G}}\mathbf{X}e^{t\mathbf{G}}dt,$ (3.5)
@@
which retrieves the well-known chain rule for the matrix commutator, cf. [32, Lem. 1], see also [9, (1.9)] for general self-adjoint operators.
@@ 3.1. Formulae for derivatives of special matrix functions
Now we have all the tools ready to calculate a symbolic derivative of a particular function. We remark that the task of differentiating a general matrix function was already investigated in the 1950s, see [34] and subsequent developments in [32] and references therein. Here, we shall be predominantly interested in the exponential, logarithm, trigonometric, and power functions, as these are fundamental in applications. Unlike in the scalar case, the same matrix derivative can always be written in multiple ways, which can be useful in various situations. The list of formulae in the following theorem includes those that are known in literature (see [7, X.4] in particular); however, we provide alternative (simple) proofs based on the formula (3.1) or (3.4), see Section 5.
@@
    $\frac{d\mathbf{A}^{r}}{d\mathbf{A}}$ $=r\int_{0}^{1}\big((1-s)\operatorname{\mathsf{L}}_{\mathbf{A}}+s\operatorname{\mathsf{R}}_{\mathbf{A}}\big)^{r-1}ds$ (3.19)
@@
Formula (3.8) for the derivative of the matrix exponential appears probably in every book on Lie algebras; see, for instance, [14, Th. 5.4] or [35, Sect. 1.2., Th. 5]. Its counterpart (3.9), which is in fact just a combination of (3.8) with the Hadamard lemma (cf. Lemma 3.4 below), is much less common, but sometimes appears in proofs of the Baker–Campbell–Hausdorff formula. Symmetrized identities (3.10), (3.11) can be found in [25, Sect. 4] and in [16, Th. 10.13], however encoded in a different language, using the Kronecker representation, which eventually relies on the interpretation of power series. The inverted versions (3.14)–(3.17) seem to be not stated in literature this explicitly and for an arbitrary positive definite matrix $\mathbf{A}$. It also appears that the potential strength of (3.17) in physics and PDE applications has not been accordingly utilized so far, see Section 4 below. The identity (3.7) is well known and is the same as (3.5) mentioned above. Formula (3.12) is the exact counterpart of (3.5), although it is unlikely to be found in the literature, since it is based on Definition 2.1. On the other hand, identity (3.13) can be directly interpreted in a usual way and it is indeed well known (cf. [16, (11.10)]), since it follows from the integral definition of the matrix logarithm (see [16, (11.1)]). The formula for the derivative of the matrix power (3.22) can be directly applied to matrix-valued partial differential equations with diffusion; see Section 4.2 below. A similar formula to (3.19) is derived in [8] to solve a certain matrix equation, which can also be directly treated using (3.4). Finally, regarding the formulae for the matrix hyperbolic and trigonometric functions, these are obviously direct consequences of the previous identities for the matrix exponential. It seems that the only relatable result in the literature is [16, Th. 12.4], which uses the Kronecker representation.
We shall demonstrate in Section 4 below that Theorem 3.2 is very useful in symbolic matrix computations that are then used to obtain theoretical results. On the other hand, the question of optimal numerical evaluation of the above derivatives and matrix functions in general is a very different and highly non-trivial problematic, to which there corresponds an enormous literature, see the overview [24] and references therein just for the computation of the matrix exponential. Let us at least provide the following worked-out example that should clarify how the above formulae can be evaluated concretely.
@@
from Theorem 3.2. Note again that the right-hand side is the composition of two fourth-order tensors: part $\operatorname{\mathsf{a\!c}}_{\mathbf{A}^{-1}}$ is expected to be there in the commutative case and the rest depends on the commutator. For simplicity, consider the case $d=2$.
Although we could apply Theorem 2.3 directly, we shall use a little trick which consists of writing
@@
$=\frac{1}{2}(\mathbf{A}^{-1}\mathbf{X}+\mathbf{X}\mathbf{A}^{-1})$
$\qquad+\frac{\theta(\delta)}{8\delta}\mathbf{A}^{-1}\big((\log\mathbf{A})^{2}\mathbf{X}-2(\log\mathbf{A})\mathbf{X}(\log\mathbf{A})+\mathbf{X}(\log\mathbf{A})^{2}\big)$
@@
This could be, of course, further manipulated using, e.g., the Cayley–Hamilton theorem for $\log\mathbf{A}$, however the core idea should already be clear. Note that the resulting formula (3.36) indeed only sees the matrix $\mathbf{A}$ and not its spectral decomposition.
@@
see, e.g., [35, p. 3, (b)]. Hence, applying also (2.9) gives